a) (12 cm, 0) b) (12 cm, 90°) d) (-12 cm, 180°) 12 cm, 180°) Which of the following is the vector (8 cm, 60°) in rectangular coordinates. d) 4i+43j . The velocity of a projectile at its highest point is always b Positive a Negative d Equal to the initial velocity ero cle all scalar quantities. a) time c) acceleration peed d) displacement
Answer #1

1.

If a vector is R = Rx i + Ry j

then magnitude |R| = sqrt (Rx^2 + Ry^2)

direction = arctan (Ry/Rx)

Now given that R = 12 cm i + 0 cm j

|R| = sqrt (12^2 + 0^2) = sqrt 144 = 12

Direction = arctan (0/12) = arctan (0) = 0 deg

R = (12 cm, 0 deg)

Correct option is A.

2.

(8 cm, 60 deg)

X-component is given by

Rx = R*cos theta

Y-component is given by

Ry = R*sin theta

then Rx = 8*cos 60 deg = 4 cm

Ry = 8*sin 60 deg = 4*sqrt 3 cm

R = 4 i + 4\sqrt 3 j

Correct option is D.

3.

Velocity of projectile at the highest point will always be zero.

Correct option is C.

4.

Scalar quantities are which have no direction,

these are: time, & speed

Correct option as A and B